Homework: Chapter 5 Homework Score: 1 of 2 47 m HW Score: 10 Problem 5-49 (algorithmic) O The doma

In this​ problem, we consider replacing an existing electrical water heater with an array of solar panels. The net installed investment cost of the panels is ​$1,235 ​($1,900 less a 35​% tax credit from the​ government). Based on an energy​ audit, the existing water heater uses 210 kilowatt hours​ (kWh) of electricity per​ month, so at ​$0.14 per​ kWh, the cost of operating the water heater is ​$29.4 per month. Assuming the solar panels can save the entire cost of heating water with​ electricity, answer the following questions.

a. What is the simple payback period for the solar​ panels?

b. What is the IRR of this investment if the solar panels have a life of 11 ​years?
Answer #1

Answer a:

Initial investment = $1235

Monthly cash inflow = Monthly saving of entire cost of heating = $29.40

As cash flows are uniform:

Simple Payback Period in months = Initial investment / Monthly cash inflow = 1235 / 29.40 = 42 months

Simple Payback Period = 42 months

\Answer b:

As monthly cash inflows are uniform we can use excel formula RATE to calculate IRR.

Given:

Initial investment = $1235

Monthly cash inflow = $29.40

Life of solar panels = 11 Years = 11 * 12 = 132 months

IRR per month is required.

IRR = RATE (nper, pmt, pv, fv, type) = RATE (132, 29.40, -1235, 0, 0) = 2.26%

Monthly IRR = 2.26%